OpenStudy (anonymous):

please Solve the equation for x: 2x + y = 6.

OpenStudy (anonymous):

x= (6+y)/2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

please Solve the equation 3x + y = 4 for y

OpenStudy (anonymous):

y=4-3x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

please Solve the equation 2a = 3a + 4c for a

OpenStudy (anonymous):

yeah you're welcome. can you give me a good answer rating?

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

theres a thing next to the reply saying good answer, just click on that